# 360-LLaMA-Factory: Plug & Play Sequence Parallelism for Long Post-Training

### v1

@@ Abstract
Haosheng Zou∗1, Xiaowei Lv∗2, Shousheng Jia1, Xiangzheng Zhang1 1Qiyuan Tech  2Renmin University zouhaosheng@163.com co-first authors
@@ 2.1 Sequence Parallelsim
2.2 Fine-Tuning framework
@@
3 Sequential Parallel Development
@@
3.1 Sequential Parallel Initialization
The initialization of sequential parallelism includes grouping of corresponding GPUs and attention replacement. Given the sequence parallel size $sp$ ($sp>1$) and gpu nums $N$, each group contains all $N/sp$ GPUs, which subsequently support communication between GPUs.
@@ 3.1 Sequential Parallel Initialization
3.2 Sequentially Parallel Data Processing
@@ 3.3 Correct Loss Calculation
where $\theta$ is the parameter of the model, $x_{i}$ is the $i_{th}$ token in the sequence, and $inst$ is the human instructions.
@@ 4 Dummy Head Ulysses
Suppose the input sequence before attention is $[bs,seq\_len/sp,hs,dim]$, where $bs$ is the batch size, $seq\_len$ is the input sequence length, $sp$ is the sequence parallel size, $hs$ is the head num, and $dim$ is hidden_dim after multi-head dimensional transformation. After the all-to-all operation, it will be converted to $[bs,seq\_len,hs/sp,dim]$. When $hs$ is not divisible by sp, the above operation will fail.
Xtuner Contributors (2023) addresses this issue by transforming the hidden dimension to introduce additional virtual heads, making it divisible. This approach involves a dimensional transformation, converting the shape from $[bs,seq\_len/sp,hs,dim]$ to $[bs,seq\_len/sp,insp\times hs,dim/insp]$, thereby expanding the number of virtual heads. Here, $insp$ represents an additional internal communication sequence parallel group, which aggregates the hidden dimension through an internal $all\_gather$ operation. This transformation reshapes the sequence into $[bs,seq\_len,hs\times insp,dim]$, followed by additional attention operations. As a result, the hidden dimension is effectively recalculated $insp$ times during the actual computation. However, in practice, this approach incurs higher memory consumption and increased communication overhead. To solve this problem, we use a simple method to add a few empty heads to solve this problem.
@@ 4.1 Dummy Head Implementation
If the number of attention heads is not divisible by the sequence parallel size $sp$, we pad the head dimension by adding $sp-(hs\%sp)$ additional heads, so that the total number of heads becomes divisible by $sp$. To ensure correctness during both the forward and backward passes, we extend the input along the head dimension before the all-to-all operation, resulting in an input shape of $[bs,seq\_len/sp,hs_{new},dim]$, where $hs_{new}$ includes the padded heads. The extra heads are then removed appropriately during the backward pass. The corresponding code is shown in Algorithm 1:
@@ 4.2 Communication Analysis
We conducted communication complexity analysis on DeepSpeed-Ulysses Jacobs et al. (2023), Ring-Attention Liu et al. (2023), USP Fang and Zhao (2024), Xtuner-Ulysses Contributors (2023), as well as our Dummy-Head-Ulysses approach. We uniformly set the batch size of the input data to $bs$, the original sequence length to $seq\_len$, and the hidden dimension to $d$, i.e., the initial input is $(bs,seq\_len,d)$. Our communication analysis accounts for the additional overhead introduced by sequence parallelism during both forward and backward propagation. The communication and time complexity comparison of all sequential parallel algorithms is shown in Table 1.
Table 1: Complexity Analysis of Different Sequence Parallel Methods
@@
Xtuner-Ulysses. Xtuner-Ulysses’ communication overhead originates from two components: all-gather and all-to-all operations. The all-to-all communication complexity remains identical to DeepSpeed-Ulysses, while an additional all-gather operation is required for query, key and value after all-to-all, contributing $O(\frac{3}{insp}\times bs\times seq\_len\times d)$. Thus, the total communication complexity is $O((\frac{8}{N}+\frac{3}{insp})\times bs\times seq\_len\times d)$, where insp represents the inner sequence parallel size. It is worth mentioning that, unlike the previous sequence parallelism, due to the operation of all_gather, the attention calculation complexity has doubled by the $insp$ times, which is caused by the increase of dim dimension.
@@
5 Thorough Analysis of Sequence Parallelism
@@ 5.1 Distributed Communication Problem
Algorithm 2: Sample code to verify distributed communication
@@
Since all-reduce is used in sequence parallelism to aggregate the final loss, it is important to note that, in actual implementation, communication should be performed using $torch.distributed.nn.all\_reduce$ rather than $torch.distributed.all\_reduce$. This is because the latter does not implement the corresponding backpropagation wrapper, and the difference between the two can be found in  [^4] https://github.com/pytorch/pytorch/issues/58005. We provide an additional code to analyze the difference between the two, as shown in Algorithm 2.
Based on a simulated sequence parallel scenario using the test code above, we observe that the gradients of $w_{0}$ under $torch.distributed.nn.all\_reduce$ are 8 and 12 (GPU 0 and GPU 1), while those under $torch.distributed.all\_reduce$ are 4 and 6. These results differ by exactly the sequence parallel size, indicating incorrect scaling in the latter case. Furthermore, when sequence parallelism is not applied (i.e., in the local setting), the gradient of $w_{0}$ is expected to be 10, which matches the averaged result between two GPUs produced by $torch.distributed.nn.all\_reduce$. This consistency further validates the correctness of using $torch.distributed.nn.all\_reduce$ for gradient aggregation in sequence parallelism.
We can perform the following analysis. Assume that the input data on two GPUs is $[x_{0},x_{1}]$, and the intermediate computation yields $[y_{0}=w_{0}x_{0},y_{1}=w_{0}x_{1}]$. The final output on each GPU becomes $[y=y_{0}+y_{1},y=y_{1}+y_{0}]$, leading to a loss of $[2y-1,2y-1]$. Under $torch.distributed.all\_reduce$, the gradient with respect to $w_{0}$ becomes $[\frac{\partial loss}{\partial y}\cdot\frac{\partial y}{\partial y_{0}}\cdot\frac{\partial y_{0}}{\partial w_{0}},\frac{\partial loss}{\partial y}\cdot\frac{\partial y}{\partial y_{1}}\cdot\frac{\partial y_{1}}{\partial w_{0}}]$. In contrast, $torch.distributed.nn.all\_reduce$ also performs an $allreduce$ operation during the backward pass, which results in a final gradient proportional to $[all\_reduce(\frac{\partial loss}{\partial y})\cdot\frac{\partial y}{\partial y_{0}}\cdot\frac{\partial y_{0}}{\partial w_{0}},all\_reduce(\frac{\partial loss}{\partial y})\cdot\frac{\partial y}{\partial y_{1}}\cdot\frac{\partial y_{1}}{\partial w_{0}}]$, which differs from the former by a factor equal to the sequence parallel size. We also evaluated the impact of this issue within our framework. As shown in Figure 1, the gradient norm differs by a factor corresponding to the sequence parallel size.
@@ 5.2 The Impact of Position IDs
It should be noted that most current decoder-only models, such as Qwen Yang et al. (2024), LLaMa Grattafiori et al. (2024), etc., use RoPE Su et al. (2024) position encoding. When using sequence parallelism, if the $position\_ids$ parameter is not explicitly passed in, the existing sequence will be re-encoded on each GPU, which will cause serious errors. That is, when no $position\_ids$ is passed in, for a sequence of length $seq\_len$, it is split into $seq\_len/sp$ length on each GPU, and it will be assigned the default $position\_ids$ of $[0,1,2,...,seq\_len/sp-1]$, Where $seq\_len$ is the length of the training data and $sp$ is the sequence parallel size. However, since it essentially represents a complete sequence, it should have a complete position encoding, that is, its position encoding should be a partition of $[0,1,2,...seq\_len-1]$. Therefore, we need to initialize the $position\_ids$ of the data in advance and explicitly pass them into the forward process of the corresponding model.
@@ 6.1 Correctness Verification
Figure 2: Comparison of SFT loss between sequential parallel and non-sequential parallel.
Figure 3: Comparison of DPO loss between sequential parallel and non-sequential parallel.
@@
Table 2: Max Sequence Length Comparison Between DeepSpeed-Ulysses (DU) and Ring-Attention (RA)
@@
Table 3: Throughout (Tokens/s) Comparsion between differnt sequence parallel methods
@@ 6.2 Performance camparsion of DeepSpeed-Ulysses and Ring-Attention
6.3 Throughout Comparsion of Different Sequence Parallel Methods

### v2

@@ Abstract
Haosheng Zou∗3, Xiaowei Lv∗2, Shousheng Jia1, Lin Li1, Xiaochun Gong1, Xiangzheng Zhang1 1Qiyuan Tech  2Renmin University  3work done at Qiyuan Tech, now at ByteDance zouhaosheng@163.com co-first authors
@@ 2.1 Sequence Parallelsim
2.2 Fine-Tuning Frameworks
@@
3 Developing Sequence Parallelism
@@
3.1 Initialization
The initialization of sequential parallelism includes grouping of corresponding GPUs and attention replacement. Given the sequence parallel size spsp ($sp>1$) and gpu nums $N$, each group contains all $N/sp$ GPUs, which subsequently support communication between GPUs.
@@ 3.1 Initialization
3.2 Data Processing
@@ 3.3 Correct Loss Calculation
where $\theta$ is the parameter of the model, $x_{i}$ is the $i_{th}$ token in the sequence, and instinst is the human instructions.
@@
3.4 SP on Vision-Language Models
To enable sequence parallelism on multimodal models, we address the challenge that visual modalities (images and videos) only have embeddings without corresponding token IDs, unlike text token IDs that can be directly partitioned. Our approach adopts a unified framework where visual content is represented using placeholder image token IDs in the input sequence, with the number of placeholders predetermined by the multimodal processor based on image characteristics.
Specifically, we construct an accompanying image position map of identical length to the input sequence, where each position contains either -1 for non-visual tokens or the corresponding visual token index (0, 1, 2, …) for image placeholders. Both the input token sequence and position map undergo identical sequence-parallel partitioning, either in a zigzag splitting pattern (zigzag-ring) or naive splitting (ulysses). During the forward pass, we employ targeted monkey patching to utilize the position map for retrieving corresponding visual embeddings from ViT outputs, effectively replacing the placeholder tokens.
This design maintains consistency with existing text-based sequence parallelism while seamlessly extending support to multimodal scenarios. Additionally, considering the significant variation in sequence lengths within multimodal batches, we implement an adaptive padding strategy that only pads sequences to the minimum required length, rather than using a fixed maximum length. This optimization reduces approximately 50% of training time on our datasets by minimizing computational overhead from excessive padding.
@@ 4 Dummy Head Ulysses
Suppose the input sequence before attention is [bs,seq_len/sp,hs,dim][bs,seq\_len/sp,hs,dim], where $bs$ is the batch size, seq_lenseq\_len is the input sequence length, $sp$ is the sequence parallel size, $hs$ is the head num, and dimdim is hidden_dim after multi-head dimensional transformation. After the all-to-all operation, it will be converted to [bs,seq_len,hs/sp,dim][bs,seq\_len,hs/sp,dim]. When $hs$ is not divisible by sp, the above operation will fail.
Xtuner Contributors (2023) addresses this issue by transforming the hidden dimension to introduce additional virtual heads, making it divisible. This approach involves a dimensional transformation, converting the shape from [bs,seq_len/sp,hs,dim][bs,seq\_len/sp,hs,dim] to $bs,seq\_len/sp,insp\times hs,dim/insp]$, thereby expanding the number of virtual heads. Here, inspinsp represents an additional internal communication sequence parallel group, which aggregates the hidden dimension through an internal all_gatherall\_gather operation. This transformation reshapes the sequence into $bs,seq\_len,hs\times insp,dim]$, followed by additional attention operations. As a result, the hidden dimension is effectively recalculated inspinsp times during the actual computation. However, in practice, this approach incurs higher memory consumption and increased communication overhead. To solve this problem, we use a simple method to add a few empty heads to solve this problem.
@@ 4.1 Dummy Head Implementation
If the number of attention heads is not divisible by the sequence parallel size $sp$, we pad the head dimension by adding $sp-(hs\%sp)$ additional heads, so that the total number of heads becomes divisible by $sp$. To ensure correctness during both the forward and backward passes, we extend the input along the head dimension before the all-to-all operation, resulting in an input shape of $bs,seq\_len/sp,hs_{new},dim]$, where $hs_{new}$ includes the padded heads. The extra heads are then removed appropriately during the backward pass. The corresponding code is shown in Algorithm 1:
@@ 4.2 Communication Analysis
We conducted communication complexity analysis on DeepSpeed-Ulysses Jacobs et al. (2023), Ring-Attention Liu et al. (2023), USP Fang and Zhao (2024), Xtuner-Ulysses Contributors (2023), as well as our Dummy-Head-Ulysses approach. We uniformly set the batch size of the input data to $bs$, the original sequence length to seq_lenseq\_len, and the hidden dimension to $d$, i.e., the initial input is (bs,seq_len,d)(bs,seq\_len,d). Our communication analysis accounts for the additional overhead introduced by sequence parallelism during both forward and backward propagation. The communication and time complexity comparison of all sequential parallel algorithms is shown in Table 1.
Table 1: Complexity analysis of different sequence parallelism methods.
@@
Xtuner-Ulysses. Xtuner-Ulysses’ communication overhead originates from two components: all-gather and all-to-all operations. The all-to-all communication complexity remains identical to DeepSpeed-Ulysses, while an additional all-gather operation is required for query, key and value after all-to-all, contributing $O(\frac{3}{insp}\times bs\times seq\_len\times d)$. Thus, the total communication complexity is $O((\frac{8}{N}+\frac{3}{insp})\times bs\times seq\_len\times d)$, where insp represents the inner sequence parallel size. It is worth mentioning that, unlike the previous sequence parallelism, due to the operation of all_gather, the attention calculation complexity has doubled by the inspinsp times, which is caused by the increase of dim dimension.
@@
5 Analysis of Sequence Parallelism
@@ 5.1 Distributed Communication Problem
Algorithm 2: Sample code to verify distributed communication.
@@
Since all-reduce is used in sequence parallelism to aggregate the final loss, it is important to note that, in actual implementation, communication should be performed using torch.distributed.nn.all_reducetorch.distributed.nn.all\_reduce rather than torch.distributed.all_reducetorch.distributed.all\_reduce. This is because the latter does not implement the corresponding backpropagation wrapper, and the difference between the two can be found in  [^4] https://github.com/pytorch/pytorch/issues/58005. We provide an additional code to analyze the difference between the two, as shown in Algorithm 2.
Based on a simulated sequence parallel scenario using the test code above, we observe that the gradients of $w_{0}$ under torch.distributed.nn.all_reducetorch.distributed.nn.all\_reduce are 8 and 12 (GPU 0 and GPU 1), while those under torch.distributed.all_reducetorch.distributed.all\_reduce are 4 and 6. These results differ by exactly the sequence parallel size, indicating incorrect scaling in the latter case. Furthermore, when sequence parallelism is not applied (i.e., in the local setting), the gradient of $w_{0}$ is expected to be 10, which matches the averaged result between two GPUs produced by torch.distributed.nn.all_reducetorch.distributed.nn.all\_reduce. This consistency further validates the correctness of using torch.distributed.nn.all_reducetorch.distributed.nn.all\_reduce for gradient aggregation in sequence parallelism.
We can perform the following analysis. Assume that the input data on two GPUs is $x_{0},x_{1}]$, and the intermediate computation yields $y_{0}=w_{0}x_{0},y_{1}=w_{0}x_{1}]$. The final output on each GPU becomes $y=y_{0}+y_{1},y=y_{1}+y_{0}]$, leading to a loss of [2y1,2y1][2y-1,2y-1]. Under torch.distributed.all_reducetorch.distributed.all\_reduce, the gradient with respect to $w_{0}$ becomes $\frac{\partial loss}{\partial y}\cdot\frac{\partial y}{\partial y_{0}}\cdot\frac{\partial y_{0}}{\partial w_{0}},\frac{\partial loss}{\partial y}\cdot\frac{\partial y}{\partial y_{1}}\cdot\frac{\partial y_{1}}{\partial w_{0}}]$. In contrast, torch.distributed.nn.all_reducetorch.distributed.nn.all\_reduce also performs an allreduceallreduce operation during the backward pass, which results in a final gradient proportional to $all\_reduce(\frac{\partial loss}{\partial y})\cdot\frac{\partial y}{\partial y_{0}}\cdot\frac{\partial y_{0}}{\partial w_{0}},all\_reduce(\frac{\partial loss}{\partial y})\cdot\frac{\partial y}{\partial y_{1}}\cdot\frac{\partial y_{1}}{\partial w_{0}}]$, which differs from the former by a factor equal to the sequence parallel size. We also evaluated the impact of this issue within our framework. As shown in Figure 1, the gradient norm differs by a factor corresponding to the sequence parallel size.
@@ 5.2 The Impact of Position IDs
It should be noted that most current decoder-only models, such as Qwen Yang et al. (2024), LLaMa Grattafiori et al. (2024), etc., use RoPE Su et al. (2024) position encoding. When using sequence parallelism, if the position_idsposition\_ids parameter is not explicitly passed in, the existing sequence will be re-encoded on each GPU, which will cause serious errors. That is, when no position_idsposition\_ids is passed in, for a sequence of length seq_lenseq\_len, it is split into seq_len/spseq\_len/sp length on each GPU, and it will be assigned the default position_idsposition\_ids of [0,1,2,,seq_len/sp1][0,1,2,...,seq\_len/sp-1], Where seq_lenseq\_len is the length of the training data and $sp$ is the sequence parallel size. However, since it essentially represents a complete sequence, it should have a complete position encoding, that is, its position encoding should be a partition of [0,1,2,seq_len1][0,1,2,...seq\_len-1]. Therefore, we need to initialize the position_idsposition\_ids of the data in advance and explicitly pass them into the forward process of the corresponding model.
@@ 6.1 Correctness Verification
Figure 2: Comparison of SFT loss with and without SP.
Figure 3: Comparison of DPO loss with and without SP.
@@
Table 2: Max sequence length comparison between DeepSpeed-Ulysses (DU) and Ring-Attention (RA).
@@
Table 3: Throughput (Tokens/s) Comparsion between differnt sequence parallel methods.
@@ 6.2 Performance camparsion of DeepSpeed-Ulysses and Ring-Attention
6.3 Throughput Comparsion of Different Sequence Parallelism Methods